Solve the following system of equations.
x+4y=5 (1)
x=6-4y (2)
What is the solution of the system?

Answer by checkley77(12844)

THERE IS NO SOLUTION BECAUSE SUBTRACTING ONE FROM THE OTHER ELIMINATES BOTH VARIABLES.
x+4y=5
x=6-4y OR X+4Y=6
X+4Y=5
-(X+4Y=6)
-----------
0X+0Y=-1 NO SOLUTION FOR X OR Y.